### v1

@@ 6.2.2 Learning methods
Gündel et al. [71] proposed a novel training strategy to handle label noise in chest radiography abnormality classification. They addressed the issue of label noise originating from natural language processed medical reports. Prior label probabilities were measured on a subset of training data re-read by 4 board-certified radiologists and were used during training to enhance the robustness of the model to label noise. These probabilities were used to adjust the weights in the loss function, making it more robust to inaccuracies in the original labels.
This is done by computing the sensitivity and specificity of the original dataset labels based on the subset of the re-read labels. Specifically, the sensitivity, $s_{sens}$ is computed as follows $s_{sens}=TP/P$ where $TP$ is the number of original positive labels which are correctly labeled based on the re-read subset and $P$ stand for the total number of positive samples in the re-read samples. $s_{spec}$ denoting specificity score $s_{spec}=TN/N$ where $TN$ is the number of original negative labels which are correctly labeled based on the re-read subset and $N$ stand for the total number of negative samples in the re-read samples.
To increase the robustness of the model, a regularization term $\mathcal{L}_{noise}$ is added to the original binary cross-entropy loss function during training.
@@
$I_{P}$ and $I_{N}$ denote the individual regularization weights for positive and negative examples, where $I_{P}^{(i)}=1-s_{sens}^{i}$ and $I_{P}^{(i)}=1-s_{spec}^{i}$. $w^{(i)}_{P}$ and $w^{(i)}_{N}$ are additional weight constants to address the problem of imbalance. $w^{(i)}_{P}=\frac{P^{(i)}+N^{(i)}}{P^{(i)}}$ and $w^{(i)}_{N}=\frac{P^{(i)}+N^{(i)}}{N^{(i)}}$, $\lambda_{Noise}$ is a weight parameter defining the overall influence of the regularization term.
In addition to the $\mathcal{L}_{noise}$, the authors proposed to exploit the correlation between labels observed in chest radiography and incorporate this information in the original loss function to further reduce the impact of label noise.
Qiu et al. [148] incorporated a regularization loss in their self-training framework, called Pathin-NL. In particular, their approach regularizes the classification model by using the KL divergence to enforce the similarity between the soft label distribution, estimated using the softmax output of the model’s current predictions, and the estimated noise free label distribution, which was computed using the softmax of the model’s output for the previous iteration. In addition, it was assumed that even in the presence of noisy labels, the majority of images were initially correctly labeled. Therefore, the original unmodified labels were also used for training their model using the standard cross-entropy loss. This will enforce the estimated label distribution not to be completely different from the initial noisy labels. They validated their approach for pathology image classification, using glioma and lung cancer datasets from The Cancer Genome Atlas (TCGA). They modeled this task as a noise label problem. In fact, each WSI is divided into multiple patches and each patch is assigned to the WSI label. Therefore, even if the WSI is labeled to be a tumor WSI, it still contains some patches of normal tissue, which are considered to be noisy samples. The results showed that the proposed method achieved an AUC of 0.872 and 0.977 on these two datasets respectively.
@@
Xiang et al. [210] proposed a weakly supervised model Graph Convolution Network-Multiple instance learning (GCN-MIL) for prostate cancer grading. It consists of 1) a self-supervised CNN model to conduct feature extraction, 2) a graph convolution network and attention pooling model to aggregate features. For self-supervised learning, they used contrastive loss on unlabeled patches (histopathological images) extracted from WSIs. For the second phase, a graph is developed based on the embedding vectors and their spatial position in the slide. DeepGCN convolution is conducted on the graph-structure data to pass information among nodes. To output a final grading prediction, the authors used attention pooling over all nodes. Since the labels are imperfect, the training process iteratively filters out noisy labels using uncertainty estimation. Noisy samples are identified based on the assumption that noisy labels samples have a large loss and uncertainty compared with clean samples. At each iteration, only clean samples within the uncertainty boundary are selected to update the GCN-MIL model.
As described in Section 6.2.2, Shi et al. [175] proposed a novel robust semi-supervised convolutional neural network Graph Temporal Ensembling (GTE). In particular, graph-based approach is used to map labeled samples of each class into a cluster, it is shown to be more beneficial than feature consistency which aims to form consensus predictions of feature representations. In contrast, feature consistency has shown significant improvement for combating noisy labels.
Yu et al. [226] presented a novel framework for pathological cancer grading. Their method takes into account space noise and level noise which tackles annotating inaccurate scope of cancerous area and cancerous level, respectively. A global-to-local rectifying strategy is used to solve the space noise. First, the ultra-large image is converted into a Multilayer Superpixel (MS) graph for data compressing. This is done using a space-aware branch which reduces the size while preserving the global features. After generating the multilayer superpixel graphs, a Graph Convolution Network (GCN) is used for preliminary classification into binary labels (positive and negative) and for generating $K$ pseudo masks by MS graphs. These graphs are used to cover category-independent regions. They are inputted into a CNN network in order to fine-tune classification results. A level-aware branch adopts different grouped kernels and a novel grading loss function to handle level noise. Finally, a bidirection co-operation are conducted to synthesize the information of the aforementioned two denoising branches which permits the assistance of the space-aware branch in grading and the feedback from the level-aware branch.
@@
Do et al. [49] presented a new Multiple Meta-model Quantifying (MMQ) [^28] https://github.com/aioz-ai/MICCAI21_MMQ method that effectively learns meta-annotation and leverages meaningful features to the medical Visual Question Answering (VQA) task. Their framework consists of three modules : 1) Meta-training for training a meta-model to extract image features used in medical VQA, 2) Data refinement which uses auto-annotation to increase the training data and deals with the noisy label by using the uncertainty of predicted score, and 3) Meta-quantifying for selecting meta-models whose robust to each others and have high accuracy during inference phase of model-agnostic tasks.
For doing meta-training, they followed Model-Agnostic Meta-Learning (MAML) [57]. Considering a model $f$ with its $\theta$ parameters. When adapting to a new classification task $T_{i}$ and its associated dataset $\{D_{i}^{tr},D_{i}^{val}\}$, the updated parameter vector ($\theta^{\prime}$) can be written as follows:
@@
where $\eta$ is a learning rate. The model parameters are trained by optimizing for the performance of $f_{\theta_{i}^{\prime}}$ with respect to $\theta$ across all tasks. At the end of each iteration, the meta-model parameters are updated using validation sets of all tasks to learn the generalized features. Formally, the meta-objective is as follows:
@@
Once the meta-training phase is done, the weights of the meta-models are used for refining the dataset. The goal of this data refinement is to update (increase) the meta-data for meta-training and remove samples that are suspected to be noisy labels or hard-to-learn. This is done by comparing the predicted scores to a predefined uncertainty threshold. If the predicted score is lower than threshold, the sample is considered noisy.
In order to identify candidate meta-models that are useful for the medical VQA task, they proposed a meta-quantifying phase. In this phase, they compute a fuse score $S_{F}$ which quantify the performance during the validating process for each meta-model :
$S_{F}=\gamma S_{P}+(1-\gamma)\sum_{t=1}^{m}1-Cosine(z_{c}^{f},z_{t}^{f})\quad\forall z_{c}^{f}\neq z_{t}^{f}$ (36)
where $S_{F}$ stands for the fuse score, $\gamma$ is the effectiveness-robustness balancing hyperparameter. $S_{P}$ is the predicted score over ground-truth label. The summation runs over all $m$ models, where $t$ is the index of the model. $z_{c}^{f}$ is the feature extracted from the current meta-model. $m$ denotes the number of candidate meta-models. $z_{t}^{f}$ is the feature extracted from the $t$-th model of the list of meta-models. Cosine represents the cosine similarity function used for checking the similarity between two features.
@@ 6.2.3 Collaborative methods
Zhou et al. [240] addressed the challenges of concept shift and noisy labels in medical imaging datasets using co-training. With the aim to tune a single target network for disease classification, they proposed to pre-train multiple reference networks to account of the uncertainty and inconsistency of the original labels. To co-optimize the target network and reduce the influence of label uncertainty, they introduced a Disentangled Distribution Learning (DDL) strategy which disentangle the multiple reference models predictions into a hard Majority Confident Label (MCL) vector (a pseudo cleaned ground-truth) and a soft Description Degree Score (DDS) vector. The MCL vector was computed by counting the number of networks giving positive and negative predictions for the corresponding disease label. The DDS vector was computed using the average over all the predictions. To optimize the target network, they employed the KL divergence based on the confidence-weighted relative entropy of the hard majority label vector with respect to the predictions of the target network. Moreover, to enforce the target network to provide consistent predictions for images with similar medical findings, the authors proposed inter- and intra-instance consistency regularization. To this end, $K$-nearest neighbor smoothing modules and image augmentation were proposed. $K$ nearest neighbors of an image (the anchor image) were computed based on the fixed soft label distribution. Then, the target network was constrained to produce similar predictions for the anchor image and its $K$ nearest neighbors. On the other hand, the anchor image was also augmented into different views and the target network was constrained to have the same prediction to these views. Experiments performed on chest X-ray and fundus image dataset, showed that the proposed approach is outperforming state-of-the-art methods.
@@
Xue et al. [217] proposed a co-training with global and local representation learning framework. Their self-ensemble co-training scheme consisted of training two independent teacher-student networks simultaneously. To ensure having different weight parameters for both networks, different image augmentation, initialization strategy were used. After one epoch of training, a Noisy Label Filter (NLF) was used to divide the data into clean and noisy samples based on the teacher encoder’s predictions. Specifically, for the NLF, a two-component Gaussian Mixture Model (GMM) is adopted. The cross-entropy loss is computed for training data, which is then inputted to the GMM. The latter fits the max-normalized loss of the training data using the Expectation-Maximization algorithm. Then, the divided clean and noisy samples are crossly sent to the peer student-teacher network as new input data. Once the noisy label detection procedure is done, rather than removing the noisy labeled sample, the authors proposed a self-supervised learning strategy.
To learn robust representation, the local contrastive loss is applied on noisy samples and the global relation loss is applied using both clean and noisy samples. For local contrastive loss, the network is encouraged to learn robust representation by creating two augmented images of the same data and regularizing the network to minimize the difference between positive pairs (two augmented views of the same source image) and maximize the difference from negative pairs (augmented images from different source). The global inter-sample relationship alignment ensures that for each batch of noisy data, the predictions of the student network and the teacher network maintain the same relationships. To this end, an inter-sample relation matrix is computed between each sample within a mini-batch. Then a global inter-sample relationship loss is added to align the inter-sample relationship of samples between the teacher and student model. Through extensive experiments on datasets such as the Histopathologic Lymph Node dataset, ISIC Melanoma dataset, the Gleason 2019 dataset, CXP dataset, the proposed approach consistently outperformed other state-of-the-art methods, especially in scenarios with high noise ratios.
@@
Zhu et al. [242] presented a robust co-teaching paradigm which cross-trains two DL networks simultaneously to select small-loss samples for training. Their proposed approach comprises two modules: 1) an Adaptive Noise Rate Estimation (ADNRE) module for estimating the dataset’s noise rate, and 2) a Consistency-based Noisy Label Correction (CNLC) module which aims to select probably clean samples based on their loss ranking and relabel highly suspected noisy samples using consistent predictions. In ADNRE, two DL networks are trained using training data and evaluated using an evaluation subset (i.e., validation subset). To estimate the noise rate of the dataset, the maximum validation accuracy is used. In CNLC, the models are initialized using the previously trained models. This module comprises two stages. In the first stage, the noisy label detection, both networks are used to select probably clean samples based on their loss ranking. In the second stage, label correction, samples with greater credible prediction are selected from suspected noisy data and relabeled, while the others are removed. To this end, at each training epoch, for estimated noisy samples, samples with the same predicted labels on both networks as well as the prediction confidences that are larger than a certain threshold, are relabeled. The idea stems from the observation that samples with consistent predictions on different models have greater credibility. Finally, the corrected samples are aggregated with small-loss samples into "a corrected set" which is used for training the network in the next iteration. The authors showed promising performance of their proposed approach to handle noisy labels, when tested on public skin lesion datasets ISIC-2017, and ISIC-20019 and a constructed thyroid ultrasound image dataset.
One drawback of co-teaching is that ordering data based on their loss may overlook difficult examples that may be correctly labeled but hard to train. In other words, cases which have similar characteristics to noisy ones but are needed for improving model performances. To overcome this issue, Peng et al. [145] proposed an improvement on co-teaching, called co-weighting.
@@
Similar to co-teaching, two DL networks are trained simultaneously, teaching each other given every mini-batch. However, in co-weighting, each network feeds forward all data and re-weight samples of the current batch dynamically. In this approach, noisy samples are identified and excluded by analyzing the statistical features of predictive history and only hard informative samples are retained. The history of prediction, i.e., all the output in the former iterations are stored in $H_{i}^{t-1}$ and is used to store the learning events that correspond to an increase of the predictions between two consecutive updates (associated with the numerical value of 1) and forgetting events that correspond to a decrease of the predictions between two consecutive updates (-1), denoted by $K_{i}^{t-1}$. It is assumed that noisily-labeled samples are forgettable in contrast clean samples undergo fewer forgettable events. Based on $K_{i}^{J}$, where $J$ is the max step of iterations, the number of forgetting events can be identified, hence the noisy samples are identified. Next, they adopted Noisy cross-validation [33] to estimate noise ratio ($\epsilon$). The reserved samples (with a ratio of 1 - $\epsilon$) underwent a ranking process.
Experiments on DigestPath2019 and the colorectal tumor dataset were conducted and showed high average accuracy ($>0.915$) of 5-fold cross-validation. The results outperformed co-teaching.
@@
Zhu et al. [241] also proposed an improvement over co-teaching framework. Similar to Peng et al. [145], the authors argued that by selecting small loss samples, co-teaching framework overlooks hard samples, which may be considered as noisy. To overcome this problem, they proposed a hard sample aware noise robust learning method. Their framework [^29] https://github.com/bupt-ai-cz/HSA-NRL/ is composed of two phases: a label correction phase and a Noise Suppressing and Hard Enhancing (NSHE) phase. The label correction phase produces an "almost clean dataset" which is then used as input in the second phase. This stage pre-discard most of the noisy samples using a self-training strategy, as described in Section 6.2.1. The NSHE phase takes the obtained "almost clean dataset" as input and outputs the final noise-robust classification model. In this phase, the authors adopted a co-learning architecture to enhance the hard samples while suppressing the few existing noisy ones. First, two DL networks, $f_{1}(x,\theta_{1})$ and $f_{2}(x,\theta_{2})$, are initialized with the same backbone and parameters. Both networks are updated using back-propagation and momentum manner. Inspired by MoCo [74], the parameters of the first DL network $\theta_{1}$ are updated by back-propagation. While the parameters of the second DL network $\theta_{2}$ are updated by:
@@
where $m\in[0,1)$ denotes a momentum coefficient. $\theta_{2}$ evolve more smoothly than $\theta_{1}$. At each epoch, $f_{2}$ selects training data for $f_{1}$ by ranking samples according to their prediction values, i.e., samples with small prediction probabilities are not used for back-propagation. The proposed method showed good performances, outperforming state-of-the-art methods across all noise ratios in tests on both medical scenario datasets (DigestPath2019 and Camelyon16) and toy computer vision datasets (CIFAR-10 and Webvision).
Co-correcting strategy [122] [^30] https://github.com/JiarunLiu/Co-Correcting, inspired also by co-teaching, aims to address the issue of noisy labels through the simultaneous training of two deep learning networks with identical architecture (also called dual-network architecture). The parameters of DL networks are updated using the principle "updated by agreement". This strategy assumes that instances with small losses are clean and collect their gradients when agreement occurs (mutual learning). The framework consists of three modules: the dual-network module based on mutual learning, the curriculum learning module, and the label updating module. The dual-network is trained by selecting clean samples based on small losses and mutual agreement. The correction of labels is based on curriculum learning, in which co-correcting introduces a label correction strategy by increasing difficulty from easy tasks to harder ones. This avoids premature label modification which modifies noise-free samples incorrectly. Finally, for the label updating module, a probabilistic estimation of whether the label is noise-free (label distribution) is done like in the PENCIL framework [223]. The idea is to update both network parameters and label estimations as label distributions. The estimated label distribution serves as a pseudo-label, it is initialized based on the noisy labels and continuously updated using backpropagation. This method was evaluated on PatchCamelyon and ISIC datasets, under different level of noise. It outperformed other state-of-the-art methods such as PENCIL, Reweight, DivideMix, Co-teaching, Co-teaching+, and Joint Optim.
@@
Knowledge distillation Knowledge distillation is a paradigm which refers to transforming knowledge from one model to the other. A student model trained from noisy labels is guided by a teacher model. Specifically, first an initial teacher model is trained on a clean dataset. In parallel, a student model is trained using a combination of original noisy labels and the output predictions of current teacher, which serves as pseudo labels to guide the training process of the student model. As the training progresses, better guidance is provided to the student model since the prediction of the teacher model becomes more reliable.

### v3

@@ 5.2.2 Learning methods
Gündel et al. [96] addressed the issue of label noise originating from natural language processed medical reports in chest radiography abnormality classification. They measured prior label probabilities on a subset of training data re-read by 4 board-certified radiologists to enhance model robustness. These probabilities were used to adjust the weights in the loss function. Sensitivity ($s_{sens}$) and specificity ($s_{spec}$) of the original dataset labels were computed based on the subset of the re-read labels. Sensitivity was defined as $s_{sens}=\frac{TP}{P}$, and specificity is $s_{spec}=\frac{TN}{N}$ where $TP$ and $TN$ are true positives and true negatives, respectively, based on the re-read subset. $P$ and $N$ are the total number of positive and negative samples in the re-read samples, respectively. To increase the robustness of the model, a regularization term $\mathcal{L}_{noise}$ was added to the binary cross-entropy loss:
@@
$I_{P}$ and $I_{N}$ are individual regularization weights for positive and negative examples, with $I_{P}^{(i)}=1-s_{sens}^{i}$ and $I_{N}^{(i)}=1-s_{spec}^{i}$. $w^{(i)}_{P}$ and $w^{(i)}_{N}$ are weight constants to address imbalance, defined as $w^{(i)}_{P}=\frac{P^{(i)}+N^{(i)}}{P^{(i)}}$ and $w^{(i)}_{N}=\frac{P^{(i)}+N^{(i)}}{N^{(i)}}$. $\lambda_{Noise}$ is a weight parameter controlling the influence of the regularization term. In addition, the authors incorporated the correlation between labels observed in chest radiography into the original loss function to further reduce the impact of label noise.
Qiu et al. [78] incorporated a regularization loss in their self-training framework, called Pathin-NL. This approach used the KL divergence to enforce the similarity between the soft label distribution, estimated using the model’s current softmax predictions, and the estimated noise free label distribution, computed using the model’s softmax output for the previous iteration. They assumed that the majority of images were initially correctly labeled. Thus, the original labels were incorporated into training via standard cross-entropy loss. This prevents the estimated label distribution from deviating significantly from the initial noisy labels. They validated their approach on pathology image classification tasks using glioma and lung cancer datasets from The Cancer Genome Atlas (TCGA). Their method achieved an AUC of 0.872 and 0.977 on the two datasets, respectively.
@@
Xiang et al. [97] proposed a weakly supervised model Graph Convolution Network-Multiple Instance Learning (GCN-MIL) for prostate cancer grading. It consists of: 1) a self-supervised CNN for feature extraction using contrastive loss on unlabeled images, 2) a GCN and attention pooling model for feature aggregation. In the second phase, a graph was constructed from embedding vectors and their spatial position. DeepGCN convolution was conducted on the graph-structure data to pass information among nodes. Attention pooling over all nodes was used for final grading prediction. To handle imperfect labels, the model iteratively filtered out noisy samples based on high loss and uncertainty, updating the GCN-MIL model with only clean samples at each iteration.
Shi et al. [95] proposed a Graph Temporal Ensembling (GTE). The graph-based approach was used to map labeled samples of each class into a cluster. It has shown to be more beneficial for semi-supervised learning than feature consistency which aims to form consensus predictions of feature representations (described in Section 5.2.2). In contrast, feature consistency has shown significant improvement for combating noisy labels.
Yu et al. [93] presented a framework for pathological cancer grading that addresses space noise (inaccurate boundaries of cancerous areas) and level noise (inaccurate cancer grading). The framework used a space-aware branch in which the large image was converted into a Multilayer Superpixel (MS) graph, significantly reducing data size while preserving the global features. These graphs were then processed with a GCN for generating pseudo-masks, which were then used by the CNN network to fine-tune the binary classification results. For handling level noise, a level-aware branch adopted grouped convolution kernels and a novel grading loss. Finally, bidirectional cooperation between both branches were conducted, achieving high performances on CAMELYON16, PANDA and HCC datasets with accuracies of 0.9472, 0.7902 and 0.5799, respectively.
@@
Do et al. [98] presented a new Multiple Meta-model Quantifying (MMQ) [^29] https://github.com/aioz-ai/MICCAI21_MMQ method designed to enhance medical Visual Question Answering (VQA) by learning meta-annotations and leveraging meaningful features. Their framework includes three modules: 1) Meta-training for training a meta-model to extract image features for medical VQA, 2) Data refinement which uses auto-annotation to increase training data and manages noisy label by evaluating the uncertainty of predicted score, and 3) Meta-quantifying for selecting meta-models whose robust to each others and have high accuracy during the inference phase of model-agnostic tasks.
For meta-training, they followed Model-Agnostic Meta-Learning (MAML) [156]. Considering a model $f$ with its parameters $\theta$, the updated parameter vector $\theta^{\prime}$ for a new task $T_{i}$ with dataset $\{D_{i}^{tr},D_{i}^{val}\}$ is given by:
@@
where $\eta$ is a learning rate. The model parameters are trained by optimizing for the performance of $f_{\theta_{i}^{\prime}}$ with respect to $\theta$ across all tasks. At the end of each iteration, the meta-model parameters are updated using validation sets of all tasks to learn generalized features. Formally, the meta-objective is as follows:
@@
After meta-training, the meta-models weights are used for data refinement, which aims to enhance the meta-data by removing samples with predicted scores below a predefined uncertainty threshold, indicating noisy samples.
The meta-quantifying phase identifies useful meta-models for the medical VQA task by computing a fuse score $S_{F}$ to quantify performance during the validating process for each meta-model :
$S_{F}=\gamma S_{P}+(1-\gamma)\sum_{t=1}^{k}1-Cosine(z_{c}^{f},z_{t}^{f})\quad\forall z_{c}^{f}\neq z_{t}^{f}$ (21)
where $S_{F}$ is the fuse score, $\gamma$ is the effectiveness-robustness balancing hyperparameter, $S_{P}$ is the predicted score over the ground-truth label, and $k$ is the number of candidate meta-models. $z_{c}^{f}$ and $z_{t}^{f}$ are the feature extracted from the current and the $t$-th meta-model, respectively. Cosine represents the cosine similarity function.
@@ 5.2.3 Collaborative methods
Zhou et al. [99] proposed a co-training approach to tune a single target network for disease classification. They pre-trained multiple reference networks to handle label uncertainty. To co-optimize the target network, they introduced a Disentangled Distribution Learning (DDL) strategy, which disentangle the multiple reference models’ predictions into a hard Majority Confident Label (MCL) vector (a pseudo cleaned ground-truth) and a soft Description Degree Score (DDS) vector. The MCL vector was computed by counting the number of networks giving positive and negative predictions for the corresponding disease label. The DDS vector was computed using the average over all the predictions. To optimize the target network, they used KL divergence based on the confidence-weighted relative entropy of the hard majority label vector with respect to the predictions of the target network. Moreover, they proposed inter- and intra-instance consistency regularization to enforce the target network to provide consistent predictions for images with similar medical findings. This involved using KNN smoothing modules and image augmentation. $K$ nearest neighbors of an image (the anchor image) were computed based on the fixed soft label distribution. Then, the target network was constrained to produce similar predictions for the anchor image and its $K$ nearest neighbors. In addition, the anchor image was also augmented into different views and the target network was constrained to have the same prediction for these views. Experiments performed on chest X-ray and fundus image dataset, showed that the proposed approach is outperforming state-of-the-art methods.
@@
Xue et al. [100] proposed a co-training with global and local representation learning framework. Two independent teacher-student networks were trained with different image augmentation and initialization strategies to ensure distinct weight parameters. After one epoch of training, a Noisy Label Filter (NLF) divided the data into clean and noisy samples based on the teacher encoder’s predictions. Specifically, the NLF used a two-component Gaussian Mixture Model (GMM) to fit the max-normalized cross-entropy loss of the training data via the Expectation-Maximization algorithm. Clean and noisy samples were then crossly sent to the peer networks. Rather than removing the noisy labeled sample, a self-supervised learning strategy was proposed. A local contrastive loss was applied on noisy samples, encouraging the network to learn robust representations by minimizing differences between augmented views of the same image and maximizing differences from other images. A global relation loss was applied to align the inter-sample relationship of samples between the teacher and student model. Experiments on datasets such as Histopathologic Lymph Node, ISIC Melanoma, Gleason 2019, and CXP showed that this approach consistently outperformed other state-of-the-art methods, especially in scenarios with high noise ratios.
@@
Zhu et al. [81] presented a robust co-teaching paradigm that cross-trains two DL networks simultaneously to select small-loss samples for training. Their approach comprises two modules. First, an Adaptive Noise Rate Estimation module was employed for estimating the dataset’s noise rate by using the maximum validation accuracy from the networks. This noise rate was used to set the percentage of small-loss samples selected as probably clean samples in the subsequent module. Second, a Consistency-based Noisy Label Correction module was applied to select probably clean samples based on their loss ranking according to both networks and to relabel highly suspected noisy samples (samples with consistent predictions and high confidence) using consistent predictions. The corrected samples were aggregated with small-loss samples into “a corrected set”, which was used for training the network in the next iteration. This approach showed promising performance when tested on public skin lesion datasets (ISIC-2017, and ISIC-20019) and a constructed thyroid ultrasound image dataset.
@@
One drawback of co-teaching is that ordering data based on their loss may overlook difficult examples that may be correctly labeled but hard to train. To overcome this issue, Peng et al. [102] proposed co-weighting, which trains two DL networks simultaneously, teaching each other with every mini-batch. Unlike co-teaching, co-weighting dynamically re-weights samples of the current batch. Noisy samples are identified and excluded by analyzing the statistical features of predictive history and only hard informative samples are retained. In this approach, the prediction history stores learning events that correspond to increases in predictions between consecutive updates and forgetting events that correspond to decreases in predictions. Noisy samples, identified by frequent forgetting events, are excluded. The noise ratio was estimated using noisy cross-validation [159]. The reserved samples underwent a ranking process. Experiments on DigestPath2019 and the colorectal tumor dataset showed high average accuracy ($>0.915$) in 5-fold cross-validation, outperforming co-teaching.
@@
Zhu et al. [80] also proposed an improvement over co-teaching framework. They proposed a hard sample aware noise robust learning method [^30] https://github.com/bupt-ai-cz/HSA-NRL/, composed of two phases: a label correction phase and a Noise Suppressing and Hard Enhancing (NSHE) phase. The label correction phase produces an “almost clean dataset” by pre-discarding most of the noisy samples using a self-training strategy (described in Section 5.2.1). The almost clean dataset is then used in the NSHE phase, which enhances hard samples while suppressing the remaining noisy ones through a colearning architecture. Two DL networks, $f_{1}(x,\theta_{1})$ and $f_{2}(x,\theta_{2})$, are initialized with the same backbone and parameters. Inspired by MoCo [137], the parameters of the first DL network $\theta_{1}$ are updated by back-propagation, while the parameters of the second DL network $\theta_{2}$ are updated using a momentum-based approach:
@@
where $m\in[0,1)$ denotes a momentum coefficient. $\theta_{2}$ evolve more smoothly than $\theta_{1}$. At each epoch, $f_{2}$ selects training data for $f_{1}$ by ranking samples according to their prediction values, excluding those with small prediction probabilities from back-propagation.
Liu et al. [101] proposed a Co-correcting strategy [^31] https://github.com/JiarunLiu/Co-Correcting to address noisy labels by simultaneously training two DL networks with identical architecture. The parameters are updated using an “updated by agreement” principle, assuming that instances with small losses are clean and collecting their gradients when agreement occurs. The framework consists of three modules. A dual-network module based on mutual learning, where networks are trained by selecting clean samples based on small losses and mutual agreement. A curriculum learning module, in which co-correcting introduces a label correction strategy by increasing difficulty from easy tasks to harder ones. Finally, a label updating module based on a probabilistic estimation of whether the label is noise-free (label distribution) similar to the PENCIL framework [160]. The idea is to update both network parameters and label estimations as label distributions. The estimated label distribution serves as a pseudo-label. It is initialized based on the noisy labels and continuously updated using backpropagation.
@@
Knowledge distillation Knowledge distillation involves transferring knowledge from one model to the other. A student model trained on noisy labels is guided by a teacher model. Initially, a teacher model is trained on a clean dataset. In parallel, a student model is trained using a combination of original noisy labels and the teacher’s output predictions, which serves as pseudo labels. As training progresses, better guidance is provided to the student model since the prediction of the teacher model becomes more reliable.